Question 1 1I s-1+2+3 +4+... +1000-7 Question 2 1.8 s-1+3+9+27+81+...-? o 121 O None of the answers are correct o 2000 O 3000000 Infinity Question 3 1.81 s-끊+흡+습+. . . +i ..?,ound your final answer to two decimal places. Question4 s- 1+4+8+16+32+64+.. +2? O 22(281) O 1+2 (21) 1.81 O None of the answers are conect o 1 + 2 (2 +1)
Answer #1

1. to find the answer, we can go with the formula S= n(1+n/2–1/2)

=>S= n(n/2+1/2)

=>S= n(n+1)/2

hence, the sum of numbers from 1 to 1000 is =1000*1001/2=500500.

2. infinity is the right answer. The sequence is going with multiplication of 3 with the number. If we proceed with the same pattern, It will goes with infinity. Hence last option is the correct.

4. option 1 is correct